# Authorship Verification based on the Likelihood Ratio of Grammar Models

### v1

@@ 5.1.1 $\mathcal{C}_{\mathrm{Enron}}$
The $\mathcal{C}_{\mathrm{Enron}}$ corpus was derived from the well-known Enron Email Dataset333available at https://www.cs.cmu.edu/~enron, which has been used in many AV and AA studies, including [17, 18, 19, 27, 31, 72, 83, 112, 4], as well as in many other research areas. The Enron email dataset has gained considerable popularity in the authorship analysis community, likely due to the fact that it is (to the best of our knowledge) the only publicly available corpus of real-world emails with a guaranteed ground truth. Similar to previous AV studies, we also reduced the initial set to a smaller number of authors (in our case from 150 to 80), as only a few suitable emails were available for some of the authors. In total, $\mathcal{C}_{\mathrm{Enron}}$ consists of 362 documents, with each document representing a concatenation of several emails from the corresponding author in order to obtain a sufficient length. Since the emails in the Enron Email Dataset contain a wide range of noise, extensive preprocessing was necessary. To this end, we decided to manually preprocess all considered texts and instead create individual replacement rules for various types of noise. First, we removed URLs, email headers, greeting/closing formulas, email signatures, (phone) numbers, quotes as well as various non-letter repetitions. Next, we normalized UTF-8 symbols and, in a final step, replaced multiple consecutive spaces, line breaks and tabs with a single space. A similar preprocessing procedure was also described by Brocardo et al. [17, 18].
@@ 5.1.2 $\mathcal{C}_{\mathrm{Wiki}}$
The $\mathcal{C}_{\mathrm{Wiki}}$ corpus consists of 752 excerpts from 288 Wikipedia talk page editors taken from the Wikipedia sockpuppets dataset, which was released by Solorio et al. [101]. The original dataset contains two partitions comprising sockpuppets and non-sockpuppets cases, where for $\mathcal{C}_{\mathrm{Wiki}}$ we only considered the latter subset. In addition, we did not use the full range of authors within the non-sockpuppets cases, as there were not sufficiently long texts available for each of the authors. Regarding the selected texts, we removed Wiki markup, timestamps, URLs and other types of noise. Furthermore, we removed sentences containing many numbers, proper names and near-duplicate string fragments as well as truncated sentences.
@@ 5.1.3 $\mathcal{C}_{\mathrm{Stack}}$
The $\mathcal{C}_{\mathrm{Stack}}$ corpus consists of 567 posts from 189 users crawled from the question-and-answer (Q & A) network Stack Exchange [^4] https://stackexchange.com in 2019. The network comprises 173 Q & A communities, with each community focusing on a specific topic (e. g., linguistics, computer science, philosophy, etc.). For the compilation of $\mathcal{C}_{\mathrm{Stack}}$, which represents a cross-topic corpus, we collected questions and answers from users who were simultaneously active in the two thematically different communities Cross Validated ($\mathcal{S}_{1}$) and Academia ($\mathcal{S}_{2}$). The verification cases were constructed as follows. Each verification case consists of exactly two documents, a known and an unknown document $\mathcal{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$. Within the Y-cases, $\mathcal{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$ were taken from $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ respectively, while within the N-cases $\mathcal{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$ were taken from the same community $\mathcal{S}_{1}$. Due to the cross-topic nature of this corpus, AV methods that utilize topic-based features are more likely to provide inverse predictions (i. e., N-cases are more likely to be classified as Y and vice versa) while AV methods that use topic-agnostic features behave more robustly. Another challenge of $\mathcal{C}_{\mathrm{Stack}}$ is that the corpus is very unbalanced with respect to the length of the documents, with lengths between 3014 and 20476 characters. To avoid stylistic distortions from the outset, we only used the original posts of the respective authors for all texts in $\mathcal{C}_{\mathrm{Stack}}$, while subsequent edits by third parties (for the purpose of spelling or stylistic improvement) were excluded [^5] Stack Exchange provides metadata to show exactly which parts of the text have been edited, and also allows the original unedited post to be retrieved.. In addition, we replaced equations/formulas, numerical values (e. g., percentages and currency values), dates, references and URLs with appropriate placeholders. The reason why we did not omit the affected sentences and instead opted for this approach was that we wanted to keep as much text content per author as possible to ensure a sufficient number of authors.
@@ 5.1.4 $\mathcal{C}_{\mathrm{ACL}}$
The $\mathcal{C}_{\mathrm{ACL}}$ corpus comprises 466 excerpts from scientific papers by a total of 233 researchers, taken from the computational linguistics archive ACL Anthology [^6] https://aclanthology.org/. The corpus was constructed in such a way that there are exactly two papers for each author [^7] Based on the metadata of the papers (more precisely, the author fields within the corresponding BibTex entries), we ensured that each paper is from exactly one author., dating from different time periods. The average time span between the two documents of each author is approximately 15.6 years, while the minimum and maximum time spans are 8 and 31 years, respectively. In addition to the temporal aspect of the documents in $\mathcal{C}_{\mathrm{ACL}}$, another challenge of this corpus is the formal language, in which the use of stylistic devices (e. g., repetitions, metaphors or rhetorical questions) is more restricted in contrast to other text types such as chat logs, forum posts or product reviews. For the original papers, we have tried to limit the content of each text as much as possible to those sections that consist mainly of natural language text (e. g., abstract, introduction, discussion, conclusion or future work). To ensure that the extracted fragments are consistent with the remarks of Bevendorff et al. [11], we decided to preprocess each paper excerpt in $\mathcal{C}_{\mathrm{ACL}}$ (semi-)manually. Among other things, we removed tables, quotations and sentences that contained a large amount of non-linguistic content such as mathematical constructs as well as specific names of researchers, systems and algorithms. We also replaced (inline) mathematical expressions, references (e. g., “[12]”, “[SHL13]”), endophoric markers, numeric values and URLs with appropriate placeholders, as we did with the texts in $\mathcal{C}_{\mathrm{Stack}}$.
@@ 5.1.5 $\mathcal{C}_{\mathrm{PJ}}$
The $\mathcal{C}_{\mathrm{PJ}}$ corpus comprises a total of 738 chat logs from 260 sex offenders collected from the Perverted-Justice [^8] http://www.perverted-justice.com portal. The chat logs were obtained from older instant messaging clients (e. g., MSN, AOL and Yahoo), whereby for each conversation we ensured that only the offender’s chat lines were extracted based on the annotations provided by the portal. To maximize variability in terms of conversation content, we selected chat lines from different messaging clients as well as from different time spans (where possible). A major challenge of $\mathcal{C}_{\mathrm{PJ}}$ is that it contains texts characterized by excessive use of non-standard English and also contains many chat-specific abbreviations and other forms of noise. This in turn makes it difficult to extract syntax-based features such as POS tags from the texts, at least when using NLP models trained on texts written in standard English (e. g., news articles or scientific documents). Another challenge of $\mathcal{C}_{\mathrm{PJ}}$ is that the text fragments it contains are very short, which in turn limits stylistic variability. In contrast to $\mathcal{C}_{\mathrm{Enron}}$, we have automatically pre-processed the texts in this corpus. Among other things, we sorted out chat lines with less than 5 tokens and 10 characters or those containing usernames, timestamps, URLs and words with many repeating characters (e. g., “mmmhhhmm”, “heeheehee” and “:-*:-*:-*:-*:-*:-*:-*:-*”). To further increase the variability in the content of each chat log $X$, we made sure that all contained lines $(x_{1},x_{2},\ldots,x_{m})$ are different from each other. For this purpose, we applied a string similarity function to each pair $(x_{i},x_{j})$ with $i\neq j$. As a metric for string similarity, we chose the Levenshtein distance [^9] Regarding the fuzzy string matching, we used FuzzyWuzzy https://github.com/seatgeek/thefuzz based on a threshold of 0.25, which we considered appropriate for our purposes. All pairs of lines that exceeded this threshold were discarded with respect to the text fragments.
@@ 5.1.6 $\mathcal{C}_{\mathrm{Apric}}$
The $\mathcal{C}_{\mathrm{Apric}}$ corpus comprises 1,395 forum posts from 284 users crawled from the portal The Apricity - A European Cultural Community [^10] https://theapricity.com. To compile the $\mathcal{C}_{\mathrm{Apric}}$, we ensured that all posts within each verification case were collected from different sub-forums covering a variety of topics, including anthropology, genetics, race and society and ethno-cultural discussion. We then pre-processed each post and removed certain markup tags, URLs, forum signatures, quotations (as well as nested quotations that contained the original author’s content), and other types of noise, similar to what we did with the texts within the $\mathcal{C}_{\mathrm{Enron}}$ corpus. However, despite the pre-processing, some textual peculiarities remained in the texts, such as truncated sentences, spelling errors and missing spaces between adjacent words or between a punctuation mark and the following word. These can be seen as a challenge of $\mathcal{C}_{\mathrm{Apric}}$, especially with regard to AV methods that use features based on word lists.
@@ 5.1.7 $\mathcal{C}_{\mathrm{Trip}}$
The $\mathcal{C}_{\mathrm{Trip}}$ corpus comprises reviews from 300 bloggers extracted from the Webis TripAdvisor Corpus 2014 published by Trenkmann and Spiel [107]. As the original corpus was not intended for the purposes of AV, we reformatted the data accordingly. While inspecting the data, we found that most of the included reviews were already of reasonable quality, so we only carried out moderate pre-processing. Among other things, we adjusted spaces next to punctuation marks (e. g., “( $\rightarrow$(alight”)) and normalized many character repetitions to a maximum of three characters (e. g., “(!!!!$\rightarrow$!!!”). Besides, we excluded reviews that contained too many named entities and we discarded authors for which only one review was available.
@@ 5.1.8 $\mathcal{C}_{\mathrm{Yelp}}$
The $\mathcal{C}_{\mathrm{Yelp}}$ corpus comprises 2,400 reviews written by 400 so-called “Yelplers” extracted from the well-known Yelp dataset [^11] https://www.yelp.com/dataset. The corpus was developed with several challenges in mind. First, compared to all other corpora we have compiled, it contains the shortest texts, with each $\mathcal{D}_{\mathcal{U}}$ consisting of at most ten sentences. For this reason, the stylistic variability within these texts is correspondingly limited. Secondly, many reviews were written at different times (in several cases often about four years apart), so linguistic changes are to be expected in reviews written by the same person. Thirdly, most of the texts in the $\mathcal{C}_{\mathrm{Yelp}}$ are restaurant reviews, which focus heavily on similar topics such as food, atmosphere and taste experiences. To a lesser extent, however, other forms of reviews are also included in the corpus (e. g., about wedding or clothing stores), which are accompanied by linguistic and also stylistic diversity with respect to the authors. Fourth, some of the reviews contain a mixture of moderate and colloquial writing styles. Regarding the latter, we observed certain spelling variations in the texts (e.g. missing apostrophes within contractions), which may result in noticeable loss of features for AV methods that rely on word lists. For the review texts in $\mathcal{C}_{\mathrm{Yelp}}$, we performed the following pre-processing steps. Sentences containing capitalized words throughout were removed (e. g., “I WAITED 4 HOURS IN THE BEND AND COULDN’T GET IN!!!!!!!!”), as were sentences containing quotes and many repetitive characters (e. g., “wait for a heart attack!!!!! Incompetent staff!!!! will never come here again!!!!”). Moreover, we removed sentences that only consist of single words (e. g., “fraud. Fraud. Lies.”) as well as review-type summaries such as “Quality: 10/10; Service: 9/10; Ambience/Location: 10/10; Overall: 9.5/10”.
@@ 5.1.10 $\mathcal{C}_{\mathrm{Blog}}$
The $\mathcal{C}_{\mathrm{Blog}}$ corpus comprises 6,000 documents written by 1,500 bloggers, taken from the well-known Blog Authorship Corpus released by Schler et al. [96]. The original dataset comprises 681,288 posts of 19,320 bloggers and was collected from blogger.com in 2004. However, a large proportion of the posts were not suitable for our purposes for various reasons. For example, many posts consisted of non-English content or contained too much noise (e. g., superfluous strings, duplicate phrases, quotes, etc.) that was difficult to preprocess without losing too much text for the respective authors. In addition, the amount of text was insufficient in itself for many authors. Therefore, we only considered a smaller fraction of the original corpus. To counteract various noise types, we first sampled sentences from the original posts and ensured that each sentence is solely restricted to (case-sensitive) English letters, spaces and standard punctuation marks. From the resulting texts we then removed sentences containing repeated characters (e. g., “?!??!”, “---” and “-.-;;”) and words (e. g., “FUN!! FUN!! FUN!!” and “kekeke”), as well as sentences with many capitalized words (e. g., “i am saying that WHOEVER HAS OPTUS PLEASE TELL ME”). Besides, we discarded sentences containing URLs, gibberish strings (e. g., “rak;db’dfth6yortyi95”, “aDr0” and “w00t”) as well as truncated sentences. Finally, we concatenated the preprocessed sentences into a single document, which in turn led to a disruption of the coherence of the resulting texts. Similar to $\mathcal{C}_{\mathrm{PJ}}$, the texts in $\mathcal{C}_{\mathrm{Blog}}$ also exhibit an excessive amount of non-standard English expressions such as “w8 4” (wait for), “b4” (before), “2day” (today), “sum1’s” (someone’s) and “jnr” (junior). This can be seen as another challenge of $\mathcal{C}_{\mathrm{Blog}}$, since the spelling of such words is not always consistently adhered to with respect to the authors.

### v2

@@ 6.1.1 $\mathcal{C}_{\mathrm{Enron}}$
The $\mathcal{C}_{\mathrm{Enron}}$ corpus was derived from the well-known Enron Email Dataset [75], which has been used in many AV and AA studies, including [23, 24, 25, 33, 38, 86, 98, 136, 6], as well as in many other research areas. The Enron email dataset has gained considerable popularity in the authorship analysis community, likely due to the fact that it is (to the best of our knowledge) the only publicly available corpus of real-world emails with a guaranteed ground truth. Similar to previous AV studies, we also reduced the initial set to a smaller number of authors (in our case from 150 to 80), as only a few suitable emails were available for some of the authors. In total, $\mathcal{C}_{\mathrm{Enron}}$ consists of 362 documents, with each document representing a concatenation of several emails from the corresponding author in order to obtain a sufficient length. Since the emails in the Enron Email Dataset contain a wide range of noise, extensive preprocessing was necessary. To this end, we decided to manually preprocess all considered texts and instead create individual replacement rules for various types of noise. First, we removed URLs, email headers, greeting/closing formulas, email signatures, (phone) numbers, quotes as well as various non-letter repetitions. Next, we normalized UTF-8 symbols and, in a final step, replaced multiple consecutive spaces, line breaks and tabs with a single space. A similar preprocessing procedure was also described by Brocardo et al. [23, 24].
@@ 6.1.2 $\mathcal{C}_{\mathrm{Wiki}}$
The $\mathcal{C}_{\mathrm{Wiki}}$ corpus consists of 752 excerpts from 288 Wikipedia talk page editors taken from the Wikipedia sockpuppets dataset, released by Solorio et al. [116]. The original dataset contains two partitions comprising sockpuppets and non-sockpuppets cases, where for $\mathcal{C}_{\mathrm{Wiki}}$ we only considered the latter subset. In addition, we did not use the full range of authors within the non-sockpuppets cases, as there were not sufficiently long texts available for each of the authors. Regarding the selected texts, we removed Wiki markup, timestamps, URLs and other types of noise. Furthermore, we removed sentences containing many numbers, proper names and near-duplicate string fragments as well as truncated sentences.
@@ 6.1.3 $\mathcal{C}_{\mathrm{Stack}}$
The $\mathcal{C}_{\mathrm{Stack}}$ corpus consists of 567 posts from 189 users crawled from the question-and-answer (Q & A) network Stack Exchange [117] in 2019. The network comprises 173 Q & A communities, with each community focusing on a specific topic (e. g., linguistics, computer science, philosophy, etc.). For the compilation of $\mathcal{C}_{\mathrm{Stack}}$, which represents a cross-topic corpus, we collected questions and answers from users who were simultaneously active in the two thematically different communities Cross Validated ($\mathcal{S}_{1}$) and Academia ($\mathcal{S}_{2}$). The verification cases were constructed as follows. Each verification case consists of exactly two documents, a known and an unknown document $\mathcal{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$. Within the Y-cases, $\mathcal{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$ were taken from $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ respectively, while within the N-cases $\mathcal{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$ were taken from the same community $\mathcal{S}_{1}$. Due to the cross-topic nature of this corpus, AV methods that utilize topic-based features are more likely to provide inverse predictions (i. e., N-cases are more likely to be classified as Y and vice versa) while AV methods that use topic-agnostic features behave more robustly. Another challenge of $\mathcal{C}_{\mathrm{Stack}}$ is that the corpus is highly unbalanced in terms of document length, with lengths between 3,014 and 20,476 characters. To avoid stylistic distortions from the outset, we only used the original posts of the respective authors for all texts in $\mathcal{C}_{\mathrm{Stack}}$, while subsequent edits by third parties (for the purpose of spelling or stylistic improvement) were excluded (Stack Exchange provides metadata to show exactly which parts of the text have been edited, and also allows the original unedited post to be retrieved.). In addition, we replaced equations/formulas, numerical values (e. g., percentages and currency values), dates, references and URLs with appropriate placeholders. The reason that we did not leave out the affected sentences and instead chose this approach was that we wanted to retain as much text content per author as possible to ensure a sufficient number of authors.
@@ 6.1.4 $\mathcal{C}_{\mathrm{ACL}}$
The $\mathcal{C}_{\mathrm{ACL}}$ corpus comprises 466 excerpts from scientific papers by a total of 233 researchers, taken from the computational linguistics archive ACL Anthology [^8] https://aclanthology.org. The corpus was constructed in such a way that there are exactly two papers for each author [^9] Using the metadata of the papers (more precisely, the author fields in the corresponding BibTex entries), we ensured that each paper was written by exactly one author., dating from different time periods. The average time span between the two documents of each author is approximately 15.6 years, while the minimum and maximum time spans are 8 and 31 years, respectively. In addition to the temporal aspect of the documents in $\mathcal{C}_{\mathrm{ACL}}$, another challenge of this corpus is the formal language, in which the use of stylistic devices (e. g., repetitions, metaphors or rhetorical questions) is more restricted in contrast to other text types such as chat logs, forum posts or product reviews. For the original papers, we have tried to limit the content of each text as much as possible to those sections that consist mainly of natural language text (e. g., abstract, introduction, discussion, conclusion or future work). To ensure that the extracted fragments are consistent with the remarks of Bevendorff et al. [14], we decided to preprocess each paper excerpt in $\mathcal{C}_{\mathrm{ACL}}$ (semi-)manually. Among other things, we removed tables, quotations and sentences that contained a large amount of non-linguistic content such as mathematical constructs as well as specific names of researchers, systems and algorithms. We also replaced (inline) mathematical expressions, references (e. g., “[12]”, “[SHL13]”), endophoric markers, numeric values and URLs with appropriate placeholders, as we did with the texts in $\mathcal{C}_{\mathrm{Stack}}$.
@@ 6.1.5 $\mathcal{C}_{\mathrm{PJ}}$
The $\mathcal{C}_{\mathrm{PJ}}$ corpus comprises a total of 738 chat logs from 260 sex offenders collected from the Perverted-Justice [^10] http://www.perverted-justice.com portal. The chat logs were obtained from older instant messaging clients (e. g., MSN, AOL and Yahoo), whereby for each conversation we ensured that only the offender’s chat lines were extracted based on the annotations provided by the portal. To maximize variability in terms of conversation content, we selected chat lines from different messaging clients as well as from different time spans (where possible). A major challenge of $\mathcal{C}_{\mathrm{PJ}}$ is that it contains texts characterized by excessive use of non-standard English and also contains many chat-specific abbreviations and other forms of noise. This in turn makes it difficult to extract syntax-based features such as POS tags from the texts, at least when using NLP models trained on texts written in standard English (e. g., news articles or scientific documents). Another challenge of $\mathcal{C}_{\mathrm{PJ}}$ is that the text fragments it contains are very short, which in turn limits stylistic variability. In contrast to $\mathcal{C}_{\mathrm{Enron}}$, we have automatically pre-processed the texts in this corpus. Among other things, we removed chat lines with less than 5 tokens and 10 characters or those containing usernames, timestamps, URLs and words with many repeating characters (e. g., “mmmhhhmm”, “heeheehee” and “:-*:-*:-*:-*:-*:-*:-*:-*”). To further increase the variability in the content of each chat log $X$, we made sure that all contained lines $(x_{1},x_{2},\ldots,x_{m})$ are different from each other. For this purpose, we applied a string similarity function to each pair $(x_{i},x_{j})$ with $i\neq j$. As a metric for string similarity, we chose the Levenshtein distance (using TheFuzz library [112]) based on a threshold of 0.25, which we considered appropriate for our purposes. All pairs of lines that exceeded this threshold were discarded with respect to the text fragments.
@@ 6.1.6 $\mathcal{C}_{\mathrm{Apric}}$
The $\mathcal{C}_{\mathrm{Apric}}$ corpus comprises 1,395 forum posts from 284 users crawled from the portal The Apricity - A European Cultural Community [7]. To compile the $\mathcal{C}_{\mathrm{Apric}}$, we ensured that all posts within each verification case were collected from different sub-forums covering a variety of topics, including anthropology, genetics, race and society and ethno-cultural discussion. We then pre-processed each post and removed certain markup tags, URLs, forum signatures, quotations (as well as nested quotations that contained the original author’s content), and other types of noise, similar to what we did with the texts within the $\mathcal{C}_{\mathrm{Enron}}$ corpus. However, despite the pre-processing, some textual peculiarities remained in the texts, such as truncated sentences, spelling errors and missing spaces between adjacent words or between a punctuation mark and the following word. These can be seen as a challenge of $\mathcal{C}_{\mathrm{Apric}}$, especially with regard to AV methods that use features based on word lists.
@@ 6.1.7 $\mathcal{C}_{\mathrm{Trip}}$
The $\mathcal{C}_{\mathrm{Trip}}$ corpus consists of reviews from 300 bloggers, which were extracted from the Webis TripAdvisor Corpus 2014, compiled by Trenkmann and Spiel [125]. Given that the original corpus was not intended for AV purposes, we reformatted the data accordingly. While inspecting the data, we found that most of the included reviews were already of reasonable quality, so we only carried out moderate pre-processing. Among other things, we adjusted spaces next to punctuation marks (e. g., “( $\rightarrow$(alight”) and normalized many character repetitions to a maximum of three characters (e. g., “(!!!!$\rightarrow$!!!”). Besides, we excluded reviews that contained too many named entities and we discarded authors for which only one review was available.
@@ 6.1.8 $\mathcal{C}_{\mathrm{Yelp}}$
The $\mathcal{C}_{\mathrm{Yelp}}$ corpus comprises 2,400 reviews written by 400 so-called “Yelplers” extracted from the well-known Yelp dataset [137]. The corpus was developed with several challenges in mind. First, compared to all other corpora we have compiled, it contains the shortest texts, with each $\mathcal{D}_{\mathcal{U}}$ consisting of at most ten sentences. For this reason, the stylistic variability within these texts is correspondingly limited. Secondly, many reviews were written at different times (in several cases often about four years apart), so linguistic changes are to be expected in reviews written by the same person. Thirdly, most of the texts in the $\mathcal{C}_{\mathrm{Yelp}}$ are restaurant reviews, which focus heavily on similar topics such as food, atmosphere and taste experiences. To a lesser extent, however, other forms of reviews are also included in the corpus (e. g., about wedding or clothing stores), which are accompanied by linguistic and also stylistic diversity with respect to the authors. Fourth, some of the reviews contain a mixture of moderate and colloquial writing styles. Regarding the latter, we observed certain spelling variations in the texts (e.g. missing apostrophes within contractions), which may result in noticeable loss of features for AV methods that rely on word lists. For the review texts in $\mathcal{C}_{\mathrm{Yelp}}$, we performed the following pre-processing steps. Sentences containing capitalized words throughout were removed (e. g., “I WAITED 4 HOURS IN THE BEND AND COULDN’T GET IN!!!!!!!!”), as were sentences containing quotes and many repetitive characters (e. g., “wait for a heart attack!!!!! Incompetent staff!!!! will never come here again!!!!”). Moreover, we removed sentences that only consist of single words (e. g., “fraud. Fraud. Lies.”) as well as review-type summaries such as “Quality: 10/10; Service: 9/10;” “Ambience/Location: 10/10; Overall: 9.5/10”.
@@ 6.1.10 $\mathcal{C}_{\mathrm{Blog}}$
The $\mathcal{C}_{\mathrm{Blog}}$ corpus comprises 6,000 documents written by 1,500 bloggers, taken from the well-known Blog Authorship Corpus, released by Schler et al. [110]. The original dataset comprises 681,288 posts of 19,320 bloggers and was collected from blogger.com in 2004. However, a large proportion of the posts were not suitable for our purposes for various reasons. For example, many posts consisted of non-English content or contained too much noise (e. g., superfluous strings, duplicate phrases, quotes, etc.) that was difficult to preprocess without losing too much text for the respective authors. In addition, the amount of text was insufficient in itself for many authors. Therefore, we only considered a smaller fraction of the original corpus. To counteract various noise types, we first sampled sentences from the original posts and ensured that each sentence is solely restricted to (case-sensitive) English letters, spaces and standard punctuation marks. From the resulting texts we then removed sentences containing repeated characters (e. g., “?!??!”, “---” and “-.-;;”) and words (e. g., “FUN!! FUN!! FUN!!” and “kekeke”), as well as sentences with many capitalized words (e. g., “i am saying that WHOEVER HAS OPTUS PLEASE TELL ME”). Besides, we discarded sentences containing URLs, gibberish strings (e. g., “rak;db’dfth6yortyi95”, “aDr0” and “w00t”) as well as truncated sentences. Finally, we concatenated the preprocessed sentences into a single document, which in turn led to a disruption of the coherence of the resulting texts. Similar to $\mathcal{C}_{\mathrm{PJ}}$, the texts in $\mathcal{C}_{\mathrm{Blog}}$ also exhibit an excessive amount of non-standard English expressions such as “w8 4” (wait for), “b4” (before), “2day” (today), “sum1’s” (someone’s) and “jnr” (junior). This can be seen as another challenge of $\mathcal{C}_{\mathrm{Blog}}$, since the spelling of such words is not always consistently adhered to with respect to the authors.